# Estimating Causal Effects for Binary Outcomes Using Per-Decision Inverse Probability Weighting

### v2

@@ Abstract
Micro-randomized trials are commonly conducted for optimizing mobile health interventions such as push notifications for behavior change. In analyzing such trials, causal excursion effects are often of primary interest, and their estimation typically involves inverse probability weighting (IPW). However, in a micro-randomized trial additional treatments can often occur during the time window over which an outcome is defined, and this can greatly inflate the variance of the causal effect estimator because IPW would involve a product of numerous weights. To reduce variance and improve estimation efficiency, we propose a new estimator using a modified version of IPW, which we call “per-decision IPW”. It is applicable when the outcome is binary and can be expressed as the maximum of a series of sub-outcomes defined over sub-intervals of time. We establish the estimator’s consistency and asymptotic normality. Through simulation studies and real data applications, we demonstrate substantial efficiency improvement of the proposed estimator over existing estimators (relative efficiency up to 1.45 and sample size savings up to 31% in realistic settings). The new estimator can be used to improve the precision of primary and secondary analyses for micro-randomized trials with binary outcomes.
Keywords Causal excursion effect; Inverse probability weighting; Log relative risk; Micro-randomized trial; Per-decision EMEE; Per-decision importance weighting.
@@ 1 Introduction
Mobile health interventions that provides individual-level support has been undergoing fast growth [1]. Smartphone- or wearable-based messages and push notifications are used for promoting antiretroviral adherence, smoking cessation, and weight maintenance [2, 3, 4]. A novel experimental design, micro-randomized trial (MRT), was proposed in recent years for developing such interventions and has been increasingly used by domain scientists [5, 6, 7, 8]. “Micro-randomization” refers to the repeated randomization among treatment options (e.g., delivering a message or not) in the trial. Such randomization can occur 100s or 1000s of times within each participant. For example, in Sense2Stop, a smoking cessation MRT, randomization occurs every minute for 10 days. In HeartSteps II [9], a physical activity MRT, randomization occurs every 5 minutes for 90 days. We describe a typical MRT design in Section 2.1.
@@
As a result, MRT data consist of longitudinal treatments and outcomes. These outcomes are called “proximal outcomes”, because they are measured in near-term after each randomization and are often the direct target of the interventions. The primary and secondary analyses of an MRT focus on the causal excursion effect of the treatment on the proximal outcome and the effect moderation by baseline variables and time-varying contexts. Causal excursion effects provide information on whether the intervention is worth further investigation (through assessing the overall effect) and, if so, how one might improve the intervention (through assessing effect moderation by contextual information). We formally introduce causal excursion effect in Section 2.2.
@@
Because of the repeated randomization, multiple treatments can occur during the time window over which the proximal outcome is defined. This time window is referred to as the “proximal outcome window.” For example, in the Sense2Stop MRT, the proximal outcome is whether the individual experienced a stress episode during the 120 minutes following a randomization, whereas randomizations occur every minute. Thus, in the 120-minute proximal outcome window, there are 119 chances for additional treatments to occur. In the HeartSteps II MRT, the proximal outcome is whether the individual takes enough steps to constitute an activity bout during the 30 minutes following a randomization, whereas randomizations occur every 5 minutes. Existing estimators use inverse probability weighting (IPW) to account for the additional treatments in the proximal outcome window [10, 11]. These estimators effectively discard the data from the time points where additional treatments do occur during the proximal outcome window, and they upweight the data from the time points where additional treatments do not occur during the proximal outcome window. Such estimators can have high variance because (i) any time point where additional treatments occur in the proximal outcome window will not contribute to the estimation; and (ii) the product of the IP weights can be large for the time points where no additional treatments occur in the proximal outcome window. Consequently, a large sample may be necessary to identify promising interventions with statistical confidence.
@@
Our main contribution is a more efficient estimator for such settings—MRTs where additional treatments may occur during the proximal outcome window. We focused on MRTs with binary proximal outcomes, which are common in practice [8, 4]. For such MRTs, the current state-of-the-art estimator was developed by [11]. We proposed a more efficient estimator for the causal excursion effect and established its consistency and asymptotic normality. The relative efficiency of the proposed estimator over that of [11] in most simulation scenarios ranges between 1.08 and 1.45, which roughly translates to savings in sample size between 7% and 31%.
@@
Our estimator is applicable as long as the binary proximal outcome can be expressed as the maximum of a series of sub-outcomes defined over sub-intervals of time. This is often the case because a binary proximal outcome is often defined as whether a specific event occurred during a pre-specified time window. Examples of such events include a stress episode during a 120-minute window [3], a physical activity bout during a 30-minute window [12], or an interaction with a smartphone app during a 3-day window [13]. The key intuition for the efficiency improvement is to realize that the value of a binary proximal outcome is determined as soon as the event of interest occurs, and there is no need to wait till the end of the proximal outcome window to assign this value. Thus, in estimation one does not need to multiply together all the IP weights but only those weights before the occurrence of the event. Consequently, (i) some data points that would otherwise have been removed (due to being multiplied with a weight 0) are retained; and (ii) for the time points where no additional treatments occur in the proximal outcome window, the product of the IP weights becomes smaller (due to fewer terms in the product). This leads to the efficiency gain. We refer to this modification of IPW as “per-decision IPW”, because the inclusion of each weight is determined by whether the event has occurred by each decision point (i.e., a time point where randomization occurs). A related modification idea of the IP weights for efficiency improvement was proposed by [14] in the reinforcement learning literature for the setting with a single continuous outcome (i.e., the cumulative reward) that is the sum of sub-outcomes. Our method applies to a different setting where the outcome itself is longitudinal binary.
@@
The rest of paper is organized as follows. In Section 1.1 we introduce the Drink Less MRT, our motivating example. In Section 2 we define the causal excursion effect using the potential outcomes framework and provide identifiability results. In Section 3 we present the new estimator and the asymptotic properties. In Section 4 we present results from simulation studies on consistency and efficiency of the proposed estimator. In Section 5 we illustrate the proposed method using the Drink Less MRT. Section 6 concludes with discussion.
@@ 1.1 Data Example: Drink Less MRT
Drink Less is a smartphone app to help users reduce harmful alcohol consumption [15, 16]. An MRT was conducted (embedded as one arm in a three-arm RCT) to assess and improve the effectiveness of a push notification intervention component that targets user engagement [13]. Each of the 349 participants in the MRT was randomized everyday at 8pm for 30 days, each time with a probability of 0.6 to receive an engagement intervention in the form of a push notification and a probability of 0.4 to receive nothing. The content of each push notification was randomly selected from a pool of messages that encourage the user to keep track of their drinking behavior using the app. The researchers have identified a large causal effect on app engagement within 1 hour of the intervention delivery and a smaller effect within 24 hours of the intervention delivery [17]. In this illustration of our method, we aim to explore whether the effect wears off within 72 hours of the intervention delivery. We define the binary proximal outcome as whether the user opens the app in the subsequent 72 hours after each randomization at 8pm. During the 72-hour proximal outcome window, there are two additional occurrences of randomization (at 8pm in the next two days), and the probability of at least one additional treatment being delivered in the proximal outcome window is 0.84 ($=1-0.4^{2}$). Therefore, on average 84% of the data points would be discarded if one were to use the estimator by [11]. Our proposed method will substantially improve the estimation efficiency.
@@ 2.1 Notation and MRT Data Structure
Assume there are $n$ individuals in an MRT. Each individual is enrolled for $T$ time points at which treatments can be delivered. We will refer to these time points as decision points. In the presentation below we assume that $T$ is the same for all individuals, but our method generalizes to settings where $T$ varies across individuals. We use capital letters without subscript $i$ to denote observations of a generic individual. Denote the treatment assignment at decision point $t$ by $A_{t}$. For simplicity we assume that the treatment assignment is binary, with $A_{t}=1$ indicating treatment and 0 indicating no treatment. Denote observations collected after decision point $t-1$ and before decision point $t$ by vector $X_{t}$. Let $X_{T+1}$ denote the information observed after decision point $T$. As will be discussed shortly, the $X_{t}$’s include both covariates and outcomes. For each individual, the entire trajectory of observed information is $O=(X_{1},A_{1},X_{2},A_{2},\ldots,X_{T},A_{T},X_{T+1})$. We assume that data from different individuals are independently and identically distributed draws from an unknown distribution $P_{0}$. Unless noted otherwise, all expectations are taken with respect to $P_{0}$.
When designing the intervention (before the MRT starts), the researchers may deem it unsafe or unethical to deliver a push notification at certain decision points—for example, when the individual is driving. This is captured by the notion of “being (in)eligible to be randomized. (The same notion has been referred to as “being (un)available” in the literature [18, 19]) Formally, $X_{t}$ contains an indicator $I_{t}$, where $I_{t}=1$ if the individual is eligible to be randomized at decision point $t$ and $I_{t}=0$ otherwise. If $I_{t}$ = 0, $A_{t}=0$ deterministically.
@@
We consider MRTs where the proximal outcome takes binary value and is defined over a window of length $\Delta$ (of decision points) following each decision point; the same proximal outcome was considered by [11]. Here $\Delta\geq 1$ is a fixed positive integer. This window is the “proximal outcome window” referred to in Section 1. Formally, suppose $Y_{t,\Delta}$ is a function of $H_{t+\Delta}$; i.e., the definition of the proximal outcome for decision point $t$ can depend on anything observed up to and including $X_{t+\Delta}$. For example, if $\Delta=1$, then $Y_{t,\Delta}$ is a function of $H_{t+1}$, i.e., is observed before $A_{t+1}$. If $\Delta=2$, then $Y_{t,\Delta}$ is a function of $H_{t+2}$, in which case additional treatments (precisely $A_{t+1}$ in this case) can occur before the value of $Y_{t,\Delta}$ is observed. Figure 1 depicts for the case of $\Delta=3$ a few consecutive decision points and the variables we defined, except for the $R$-variables which will be defined in Section 2.3.
@@
In the Drink Less MRT, the decision point is everyday at 8pm. Each individual is in the study for $T=30$ days. $A_{t}$ is whether a push notification is delivered at 8pm on day $t$. $X_{t}$ include baseline variables and time-varying variables such as the number of days in the study and user’s previous interaction with the app. The researchers deemed that it is always appropriate to deliver push notifications, so $I_{t}=1$ for all $t$. The proximal outcome, $Y_{t,\Delta}$, is whether the individual logged in to the app in the subsequent 72 hours; thus $\Delta=3$.
@@ 2.2 Potential Outcomes and Causal Excursion Effects
For a mobile health intervention, causal quantities of interest include whether the intervention has an impact on an individual’s behavior and how this impact varies with contextual information such as the number of days in the study. To define the causal effects precisely we use the potential outcomes framework [20, 21]. Lower case letters represent instantiations (non-random values) of the corresponding capital letter random variable. For example, $a_{t}$ is an instantiation of treatment $A_{t}$. For every individual, denote by $X_{t}(\bar{a}_{t-1})$ the $X_{t}$ that would have been observed at decision point $t$ if the individual had been assigned a treatment sequence of $\bar{a}_{t-1}$ prior to $t$. The potential outcome of $H_{t}$ under $\bar{a}_{t-1}$ is
@@
We consider the causal excursion effect (CEE) of $A_{t}$ on $Y_{t,\Delta}$ on the log relative risk scale [11], defined as
@@
Here $\bar{0}_{\Delta-1}$ is a vector of 0’s of length $\Delta-1$; $(\bar{A}_{t-1},1,\bar{0}_{\Delta-1})$ and $(\bar{A}_{t-1},0,\bar{0}_{\Delta-1})$ are sequences of treatment assignments of length $t+\Delta-1$. $S_{t}(\bar{A}_{t-1})$ is a length-$p$ vector of summary variables from $H_{t}(\bar{A}_{t-1})$, usually chosen by the domain scientist based on the scientific question regarding effect modification. In Drink Less MRT, we will first set $S_{t}(\bar{A}_{t-1})=1$ to obtain the fully marginal effect (i.e., averaged over all participants and all days in the study) of a push notification on the app use in the subsequent 72-hour window. We will then set $S_{t}(\bar{A}_{t-1})=(1,\text{Day}_{t})$ to assess how the effect is moderated by days in the study.
@@
The causal excursion effect (2) is commonly used in the primary and secondary analyses for MRTs [22, 4]. It differs from most causal effects under time-varying treatments such as marginal structural models and structural nested mean models [23, 24]. Instead of contrasting fixed treatment trajectories, (2) contrasts between two stochastic treatment trajectories, $(\bar{A}_{t-1},1,\bar{0}_{\Delta-1})$ and $(\bar{A}_{t-1},0,\bar{0}_{\Delta-1})$. That is, (2) is a causal contrast between two excursions from the treatment policy: following the treatment policy up to time $t-1$ and then deviating from the policy to assign $A_{t}=1$ and no treatment for the next $\Delta-1$ decision points, and following the treatment policy up to time $t-1$ and then deviating to assign $A_{t}=0$ and no treatment for the next $\Delta-1$ decision points. (2) is conditional on $I_{t}(\bar{A}_{t-1})=1$ and $S_{t}(\bar{A}_{t-1})$, meaning that it concerns the subpopulation who, after following the treatment policy up to time $t-1$, are eligible to be randomized at $t$ and within a stratum defined by $S_{t}(\bar{A}_{t-1})$. In practice we often impose a model on how $\text{CEE}\left\{S_{t}(\bar{A}_{t-1})\right\}$ depends on $S_{t}(\bar{A}_{t-1})$ to pool across $S_{t}$-strata.
@@
$\text{CEE}\left\{S_{t}(\bar{A}_{t-1})\right\}$ depends on the treatment policy in the MRT, because the variables in $H_{t}(\bar{A}_{t-1})\setminus S_{t}(\bar{A}_{t-1})$ are marginalized over in the conditional expectations in (2). Such dependence is scientifically desirable for two reasons. First, the causal excursion effect approximates the treatment effect in a real-world implementation. A good MRT policy would already have incorporated implementation considerations such as burden and feasibility through the choice of randomization probability and availability criteria, so a feasible policy will not deviate too far from the MRT policy. Second, the causal excursion effect indicates the effective deviations from the MRT policy and how the policy might be improved. A fully marginal effect (by setting $S_{t}(\bar{A}_{t-1})=1$ in (2)) indicates whether a treatment is worth further investigation, and an effect modification analysis (by setting $S_{t}(\bar{A}_{t-1})$ to be certain time-varying covariates in (2)) indicates whether the MRT policy should be modified to depend on time-varying covariates. In addition, the dependence of $\text{CEE}\left\{S_{t}(\bar{A}_{t-1})\right\}$ on the MRT policy resembles the primary analysis in factorial designs and allows one to design trials with a higher power to detect meaningful effects. Related marginalization ideas were considered by [25, 26]. See [18], [11], and [19] for more discussion on causal excursion effect and its use in MRT. See [27] for a comprehensive comparison of various causal estimands under time-varying treatments.

### v3

@@ Abstract
Micro-randomized trials are commonly conducted for optimizing mobile health interventions such as push notifications for behavior change. In analyzing such trials, causal excursion effects are often of primary interest, and their estimation typically involves inverse probability weighting (IPW). However, in a micro-randomized trial, additional treatments can often occur during the time window over which an outcome is defined, and this can greatly inflate the variance of the causal effect estimator because IPW would involve a product of numerous weights. To reduce variance and improve estimation efficiency, we propose two new estimators using a modified version of IPW, which we call “per-decision IPW”. The second estimator further improves efficiency using the projection idea from the semiparametric efficiency theory. These estimators are applicable when the outcome is binary and can be expressed as the maximum of a series of sub-outcomes defined over sub-intervals of time. We establish the estimators’ consistency and asymptotic normality. Through simulation studies and real data applications, we demonstrate substantial efficiency improvement of the proposed estimator over existing estimators. The new estimators can be used to improve the precision of primary and secondary analyses for micro-randomized trials with binary outcomes.
@@ 1 Introduction
Mobile health interventions that provide individual-level support have been undergoing fast growth (Silva and others, 2015). Smartphone- or wearable-based messages and push notifications are used for promoting antiretroviral adherence, smoking cessation, and weight maintenance (Free and others, 2013; Battalio and others, 2021; Klasnja and others, 2021). A novel experimental design, micro-randomized trial (MRT), was proposed in recent years for developing such interventions and has been increasingly used by domain scientists (Dempsey and others, 2015; Liao and others, 2016; Klasnja and others, 2015; Rabbi and others, 2018). “Micro-randomization” refers to the repeated randomization among treatment options (e.g., delivering a message or not) in the trial. Such randomization can occur 100s or 1000s of times within each participant. For example, in Sense2Stop, a smoking cessation MRT, randomization occurs every minute for 10 days. In HeartSteps II (Spruijt-Metz and others, 2022), a physical activity MRT, randomization occurs every 5 minutes for 90 days. We describe a typical MRT design in Section 2.1.
@@
As a result, MRT data consist of longitudinal treatments and outcomes. These outcomes are called “proximal outcomes”, because they are measured in the near term after each randomization and are often the direct target of the interventions. The primary and secondary analyses of an MRT focus on the causal excursion effect of the treatment on the proximal outcome and the effect moderation by baseline variables and time-varying contexts. Causal excursion effects provide information on whether the intervention is worth further investigation (through assessing the fully marginal effect) and, if so, how one might improve the intervention (through assessing effect moderation by contextual information). We formally introduce the causal excursion effect in Section 2.2.
@@
Because of the repeated randomization, multiple treatments can occur during the time window over which the proximal outcome is defined. This time window is referred to as the “proximal outcome window.” For example, in the Sense2Stop MRT, the proximal outcome is whether the individual experienced a stress episode during the 120 minutes following a randomization, whereas randomizations occur every minute. Thus, in the 120-minute proximal outcome window, there are 119 chances for additional treatments to occur. In the HeartSteps II MRT, the proximal outcome is whether the individual takes enough steps to constitute an activity bout during the 30 minutes following a randomization, whereas randomizations occur every 5 minutes. Existing estimators use inverse probability weighting (IPW) to account for the additional treatments in the proximal outcome window (Dempsey and others, 2020; Qian and others, 2021). These estimators effectively discard the data from the time points where additional treatments do occur during the proximal outcome window, and they upweight the data from the time points where additional treatments do not occur during the proximal outcome window. Such estimators can have high variance because (i) any time point where additional treatments occur in the proximal outcome window will not contribute to the estimation; (ii) the product of the IP weights can be large for the time points where no additional treatments occur in the proximal outcome window. Consequently, a large sample may be necessary to identify promising interventions with statistical confidence.
@@
Our main contribution is two more efficient estimators for such settings—MRTs where additional treatments may occur during the proximal outcome window. We focused on MRTs with binary proximal outcomes, which are common in practice (Rabbi and others, 2018; Klasnja and others, 2021). For such MRTs, the current state-of-the-art estimator was developed by Qian and others (2021). We proposed two more efficient estimators for the causal excursion effect and established its consistency and asymptotic normality. The relative efficiency of the proposed estimators over that of Qian and others (2021) in most simulation scenarios ranges between 1.08 and 1.45 or even higher, which roughly translates to savings in sample size between 7% and 31% or even larger. The efficiency improvement is greater when there is a higher chance of one or more additional treatments occurring in the proximal outcome window; for example, when the proximal outcome window consists of more decision points or when the randomization probability is larger. When the proximal outcome window doesn’t contain additional decision points (i.e., when the proximal outcome is measured immediately), the proposed method is equivalent to the EMEE in Qian and others (2021).
@@
Our estimators are applicable as long as the binary proximal outcome can be expressed as the maximum of a series of sub-outcomes defined over sub-intervals of time. This is often the case because a binary proximal outcome is often defined as whether a specific event occurred during a pre-specified time window. Examples of such events include a stress episode during a 120-minute window (Battalio and others, 2021), a physical activity bout during a 30-minute window (Liao and others, 2020), or an interaction with a smartphone app during a 3-day window (Bell and others, 2020). The key intuition for the efficiency improvement is to realize that the value of a binary proximal outcome is determined as soon as the event of interest occurs, and there is no need to wait till the end of the proximal outcome window to assign this value. Thus, when performing estimation, one does not need to multiply together all the IP weights but only those weights before the occurrence of the event. Consequently, (i) some data points that would otherwise have been removed (due to being multiplied with a weight 0) are retained; and (ii) for the time points where no additional treatments occur in the proximal outcome window, the product of the IP weights becomes smaller (due to fewer terms in the product). This leads to the efficiency gain. We refer to this modification of IPW as “per-decision IPW”, because the inclusion of each weight is determined by whether the event has occurred by each decision point (i.e., a time point where randomization occurs). A related modification idea of the IP weights for efficiency improvement was proposed by (Precup, 2000) in the reinforcement learning literature for the setting with a single continuous outcome (i.e., the cumulative reward) that is the sum of sub-outcomes. Our method applies to a different setting where the outcome itself is longitudinal binary. Furthermore, our second estimator further improves efficiency by leveraging the projection idea from the semiparametric efficiency theory (Bickel and others, 1993; Murphy and others, 2001).
@@
The rest of the paper is organized as follows. In Section 1.1 we introduce the Drink Less MRT, our motivating example. In Section 2 we define the causal excursion effect using the potential outcomes framework and provide identifiability results. In Section 3 we present the first new estimator, pd-EMEE, which capitalizes on the per-decision IPW idea, and the asymptotic properties of this estimator. In Section 4, we present the second new estimator, pd-EMEE2, which further improves efficiency using the projection idea from the semiparametric efficiency theory, and derive the asymptotic properties of this estimator. In Section 5 we present results from simulation studies on the consistency and efficiency of the two proposed estimators. In Section 6 we illustrate the proposed method using the Drink Less MRT and HeartSteps I MRT datasets. Section 7 concludes with a discussion.
@@ 1.1 Data Example: Drink Less MRT
Drink Less is a smartphone app to help users reduce harmful alcohol consumption (Garnett and others, 2019, 2021). An MRT was conducted (embedded as one arm in a three-arm RCT) to assess and improve the effectiveness of a push notification intervention component that targets user engagement (Bell and others, 2020). Each of the 349 participants in the MRT was randomized every day at 8 pm for 30 days, each time with a probability of 0.6 to receive an engagement intervention in the form of a push notification and a probability of 0.4 to receive nothing. The content of each push notification was randomly selected from a pool of messages that encouraged the user to keep track of their drinking behavior using the app. The researchers have identified a large causal effect on app engagement within 1 hour of the intervention delivery and a smaller effect within 24 hours of the intervention delivery (Bell and others, 2023). In this illustration of our method, we aim to explore how this effect depends on the choice of the time window, from 1 day to 5 days. We define the binary proximal outcome as whether the user opens the app in the subsequent $\Delta$ days after each randomization at 8 pm. For example, when $\Delta=3$, during the 3-day proximal outcome window there are two additional occurrences of randomization (at 8 pm in the next two days), and the probability of at least one additional treatment being delivered in the proximal outcome window is 0.84 ($=1-0.4^{2}$). Therefore, on average 84% of the data points would be discarded if one were to use the estimator by Qian and others (2021). Our proposed methods will substantially improve the estimation efficiency.
1.2 Data Example: HeartSteps I MRT
HeartSteps I MRT aimed at developing mHealth interventions for increasing physical activity among sedentary adults (Klasnja and others, 2015). In illustrating our method, we focus on one of the intervention components in the MRT, the activity suggestions. Each of the 37 participants was randomized five times a day for 42 days, each time with a probability of 0.6 to receive a contextually tailored activity suggestion in the form of a push notification, and a probability of 0.4 to receive nothing. The researchers have identified a substantial causal effect on steps taken within 30 minutes of the intervention delivery (Klasnja and others, 2018). We aim to explore whether the activity suggestion has a prolonged effect in the next 24 hours. We define the binary proximal outcome as whether the participant takes more steps than a pre-determined threshold (e.g., 8,000 steps) in the 24 hours following an intervention delivery. During the 24-hour proximal outcome window, there are four additional occurrences of randomization, and thus our proposed method can improve the estimation efficiency.
@@ 2.1 Notation and MRT Data Structure
Assume there are $n$ individuals in an MRT. Each individual is enrolled for $T$ time points at which treatments can be delivered. We will refer to these time points as decision points. In the presentation below we assume that $T$ is the same for all individuals, but our method generalizes easily to settings where $T$ varies across individuals. We use capital letters without subscript $i$ to denote observations of a generic individual. Denote the treatment assignment at decision point $t$ by $A_{t}$. For simplicity, we assume that the treatment assignment is binary, with $A_{t}=1$ indicating treatment and 0 indicating no treatment. Denote observations collected after decision point $t-1$ and before decision point $t$ by vector $X_{t}$. Let $X_{T+1}$ denote the information observed after decision point $T$. As will be discussed shortly, the $X_{t}$’s include both covariates and outcomes. For each individual, the entire trajectory of observed information is $O=(X_{1},A_{1},X_{2},A_{2},\ldots,X_{T},A_{T},X_{T+1})$. We assume that data from different individuals are independently and identically distributed draws from an unknown distribution $P_{0}$. Unless noted otherwise, all expectations are taken with respect to $P_{0}$.
When designing the intervention (before the MRT starts), the researchers may deem it unsafe or unethical to deliver a push notification at certain decision points—for example, when the individual is driving. This is captured by the notion of “being (un)available to be randomized” (Boruvka and others, 2018). Formally, $X_{t}$ contains an indicator $I_{t}$, where $I_{t}=1$ if the individual is available to be randomized at decision point $t$ and $I_{t}=0$ otherwise. If $I_{t}$ = 0, $A_{t}=0$ deterministically.
@@
We consider MRTs where the proximal outcome takes a binary value and is defined over a window of length $\Delta$ (of decision points) following each decision point; the same proximal outcome was considered by Qian and others (2021). Here $\Delta\geq 1$ is a fixed positive integer. This window is the “proximal outcome window” referred to in Section 1. Formally, suppose $Y_{t,\Delta}$ is a function of $H_{t+\Delta}$; i.e., the definition of the proximal outcome for decision point $t$ can depend on anything observed up to time $t+\Delta$ and including $X_{t+\Delta}$. For example, if $\Delta=1$, then $Y_{t,\Delta}$ is a function of $H_{t+1}$, i.e., is observed before $A_{t+1}$. If $\Delta=2$, then $Y_{t,\Delta}$ is a function of $H_{t+2}$, in which case additional treatments (precisely $A_{t+1}$ in this case) can occur before the value of $Y_{t,\Delta}$ is observed. Figure 1 depicts for the case of $\Delta=3$ a few consecutive decision points and the variables we defined, except for the $R$-variables which will be defined in Section 2.3.
@@
In the Drink Less MRT, the decision point is every day at 8 pm. Each participant is in the study for $T=30$ days. $A_{t}$ is whether a push notification is delivered at 8 pm on day $t$. $X_{t}$ includes baseline variables and time-varying variables such as the number of days in the study and the user’s previous interaction with the app. The researchers deemed that it is always appropriate to deliver push notifications, so $I_{t}=1$ for all $t$. The proximal outcome, $Y_{t,\Delta}$, is whether the participant logged in to the app in the subsequent $\Delta$ days, where we consider $\Delta=1,\ldots,5$ in separate analyses.
In the HeartSteps I MRT, the decision point occurs five times a day during the morning commute, lunchtime, mid-afternoon, evening commute, and after dinner. Each participant is in the study for $T=210$ decision points (which amounts to 42 days). $A_{t}$ is whether an activity suggestion is delivered at decision point $t$. $X_{t}$ includes baseline variables and time-varying variables such as the participant’s previous step count. $I_{t}=0$ if the participant is driving, already walking, or with no internet connection. The proximal outcome, $Y_{t,\Delta}$, is whether the participant takes more steps than a given threshold in the next 24 hours (so $\Delta=5$). We consider a few thresholds from 5,000 to 12,000 steps in separate analyses.
@@ 2.2 Potential Outcomes and Causal Excursion Effects
For a mobile health intervention, causal quantities of interest include whether the intervention has an impact on an individual’s behavior and how this impact varies with contextual information such as the number of days in the study. To define the causal effects precisely we use the potential outcomes framework (Rubin, 1974; Robins, 1986). Lowercase letters represent instantiations (non-random values) of the corresponding capital letter random variable. For example, $a_{t}$ is an instantiation of treatment $A_{t}$. For every individual, denote by $X_{t}(\bar{a}_{t-1})$ the $X_{t}$ that would have been observed at decision point $t$ if the individual had been assigned a treatment sequence of $\bar{a}_{t-1}$ prior to $t$. The potential outcome of $H_{t}$ under $\bar{a}_{t-1}$ is
@@
We consider the causal excursion effect (CEE) of $A_{t}$ on $Y_{t,\Delta}$ on the log relative risk scale (Qian and others, 2021), defined as
@@
Here $\bar{0}_{\Delta-1}$ is a vector of 0’s of length $\Delta-1$; $(\bar{A}_{t-1},1,\bar{0}_{\Delta-1})$ and $(\bar{A}_{t-1},0,\bar{0}_{\Delta-1})$ are sequences of treatment assignments of length $t+\Delta-1$. $S_{t}(\bar{A}_{t-1})$ is a length-$p$ vector of summary variables from $H_{t}(\bar{A}_{t-1})$, usually chosen by the domain scientist based on the scientific question regarding effect modification. In Drink Less MRT, we will first set $S_{t}(\bar{A}_{t-1})=1$ to obtain the fully marginal effect (i.e., averaged over all participants and all days in the study) of a push notification on the app use in the subsequent $\Delta$ days. We will then set $S_{t}(\bar{A}_{t-1})=(1,\text{Day}_{t})$ to assess how the effect is moderated by days in the study.
@@
The causal excursion effect (2) is commonly used in the primary and secondary analyses for MRTs (Nahum-Shani and others, 2021; Klasnja and others, 2021). It differs from most causal effects under time-varying treatments such as marginal structural models and structural nested mean models (Robins, 1994; Robins and others, 2000). Instead of contrasting fixed treatment trajectories, (2) contrasts between two stochastic treatment trajectories, $(\bar{A}_{t-1},1,\bar{0}_{\Delta-1})$ and $(\bar{A}_{t-1},0,\bar{0}_{\Delta-1})$. That is, (2) is a causal contrast between two excursions from the treatment policy: following the treatment policy up to time $t-1$ and then deviating from the policy to assign $A_{t}=1$ and no treatment for the next $\Delta-1$ decision points, and following the treatment policy up to time $t-1$ and then deviating to assign $A_{t}=0$ and no treatment for the next $\Delta-1$ decision points. (2) is conditional on $I_{t}(\bar{A}_{t-1})=1$ and $S_{t}(\bar{A}_{t-1})$, meaning that it concerns the subpopulation who, after following the treatment policy up to time $t-1$, are available to be randomized at $t$ and within a stratum defined by $S_{t}(\bar{A}_{t-1})$. In practice we often impose a model on how $\text{CEE}\left\{S_{t}(\bar{A}_{t-1})\right\}$ depends on $S_{t}(\bar{A}_{t-1})$ to pool across $S_{t}$-strata.
@@
$\text{CEE}\left\{S_{t}(\bar{A}_{t-1})\right\}$ depends on the treatment policy in the MRT, because the variables in $H_{t}(\bar{A}_{t-1})\setminus S_{t}(\bar{A}_{t-1})$ are marginalized over in the conditional expectations in (2). Such dependence is scientifically desirable for two reasons. First, the causal excursion effect approximates the treatment effect in a real-world implementation. A good MRT policy would already have incorporated implementation considerations such as burden and feasibility through the choice of randomization probability and availability criteria, so a feasible policy will not deviate too far from the MRT policy. Second, the causal excursion effect indicates the effective deviations from the MRT policy and how the policy might be improved. A fully marginal effect (by setting $S_{t}(\bar{A}_{t-1})=1$ in (2)) indicates whether a treatment is worth further investigation, and an effect modification analysis (by setting $S_{t}(\bar{A}_{t-1})$ to be certain time-varying covariates in (2)) indicates whether the MRT policy should be modified to depend on time-varying covariates. In addition, the dependence of $\text{CEE}\left\{S_{t}(\bar{A}_{t-1})\right\}$ on the MRT policy resembles the primary analysis in factorial designs and allows one to design trials with a higher power to detect meaningful effects. Related marginalization ideas were considered by (Robins, 2004; Neugebauer and others, 2007). See Boruvka and others (2018), Qian and others (2021), and Qian and others (2022) for more discussion on causal excursion effect and its use in MRT. See Guo and others (2021) for a comprehensive comparison of various causal estimands under time-varying treatments.